# From homeostasis to resource sharing: Biologically and economically aligned multi-objective multi-agent AI safety benchmarks

### v4

@@ Abstract
Developing safe, aligned agentic AI systems requires comprehensive empirical testing, yet many existing benchmarks neglect crucial themes aligned with biology and economics, both time-tested fundamental sciences describing our needs and preferences. To address this gap, the present work focuses on introducing biologically and economically motivated themes that have been neglected in current mainstream discussions on AI safety - namely a set of multi-objective, multi-agent alignment benchmarks that emphasize homeostasis for bounded and biological objectives, diminishing returns for unbounded, instrumental, and business objectives, sustainability principle, and resource sharing. We implemented eight main benchmark environments on the above themes, to illustrate key pitfalls and challenges in agentic AI-s, such as unboundedly maximizing a homeostatic objective, over-optimizing one objective at the expense of others, neglecting safety constraints, or depleting shared resources.
Keywords Value Alignment, Multi-Objective AI, Evaluation, Human Values, Nonlinear Utility Functions, Homeostasis, Bounded Objectives, Safety Constraints, Unbounded Objectives, Diminishing Returns, Sustainability
@@ 1 Introduction
This work introduces safety challenges for an agent’s ability to learn and act in desired ways in relation to biologically and economically relevant aspects. In total, we implemented nine benchmarks, which are conceptually divided into three developmental stages: “basic biologically inspired dynamics in objectives”, “multi-objective agents”, and “cooperation”. The first two stages can be considered as proto-cooperative stages, since the behavioral dynamics tested in these benchmarks will be later potentially very relevant for supporting and enabling cooperative behavior in multi-agent scenarios.
@@
A common theme over the environments below: At some near-future stage, the same model of the agent should be able to perform well in all of these environments and inside all different agent bodies related to these environments. This is important to avoid overfitting different models to different environments and bodies. However, it is also possible to train separate models for each environment. The environments are randomized through seeds to avoid the most basic form of misgeneralization from occurring. This contrasts to DeepMind’s AI Safety Gridworlds Leike et al. (2017), where the maps had a fixed layout.
The benchmarks are implemented in a gridworld-based environment. The gridworld-based environment is extended from and mostly compatible with DeepMind’s AI Safety Gridworlds Leike et al. (2017) code. Therefore, it is easy to adapt any existing DeepMind’s AI safety gridworlds-based environment to support multiple objectives, multiple agents, and many other new functionalities.
To support more robust multi-agent simulations, the framework is made compatible with PettingZoo Terry et al. (2020) and OpenAI Gym Brockman et al. (2016) APIs as well.
The source code for the extended gridworlds framework can be found at  [^1] https://github.com/levitation-opensource/ai-safety-gridworlds/tree/biological-compatibility-benchmarks. The source code for concrete implementation of biologically compatible benchmarks described in this publication, as well as code for running the agents can be found at  [^2] https://github.com/aintelope/biological-compatibility-benchmarks. The latter also contains an example code for a random agent.
@@ 1.1 Summary of major themes in benchmarks below
Modern AI benchmark environments often default to maximizing unbounded rewards, overlooking nuanced dynamics that would align with real-world constraints. Our suite of eight main benchmarks introduces the following:
Homeostasis and bounded objectives - Agents must maintain certain internal metrics (e.g., food / water levels) within safe ranges, penalizing both deficiency and excess Betts (2017). This inverted U-shaped reward structure prevents the "utility monster" behavior introduced in Nozick (1974) by demonstrating that too much of a good thing can be harmful even for the very same objective that was maximized for. This seems to apply to most or even all biological objectives. This means that agents need to be able to understand the dynamics of inverted U-shaped rewards, and the agents should not be greedy.
Secondly, incorporating bounded objectives also reduces the stakes of optimization processes, thus reducing the incentives for extreme actions and manifestations of Goodhart’s law.
@@
This includes balancing between multiple unbounded objectives, which applies mostly to business performance objectives. Using concepts from economics Krugman and Wells (2013), these performance objectives with "concave utility functions" can alternatively be represented by "convex indifference curves".
@@
Balancing multiple objectives can also be considered as a mitigation against Goodhart’s law. The principle of multi-objective diminishing returns and its relation to Goodhart’s law is described in more detail in Smith et al. (2022).
@@
Vamplew et al. (2022) explain the need for multiple objectives, utility functions, and including safety considerations in the plurality of objectives. Furthermore, Vamplew et al. (2021) propose a thresholded lexicographic ordering, which aims to first maximize the value of thresholded (safety) objectives, and then secondarily maximize the unthresholded (performance) objectives. In other words, the safety objectives could be treated as having hierarchically higher priority than the performance objectives, until the safety objectives have been satisfied to a specified threshold value.
Along these lines, we propose an additional perspective: The distinction between constraints versus objectives in combinatorial optimization Korte and Vygen (2006) is analogous to the distinction between safety objectives and performance objectives.
It is notable that in combinatorial optimization problems, the concept of constraints has been naturally considered as part of the setup, along with the concept of objectives. In contrast, unfortunately, this has often not been the case in the use of reinforcement learning. However, in reality, both performance and safety objectives should be present. Sutton and Barto (2018) mentions the general concept of constraints, although it is not directly related to safety. We propose that safety objectives can be often considered similar to constraints in the paradigm of combinatorial optimization, while performance objectives and rewards would be similar to the objective functions.
@@
Sustainability - The environment contains resources that renew, but if these resources are consumed too quickly, they run out and therefore cannot replenish anymore. This is especially relevant in multi-agent environments. The agent depends on these resources for long-term survival. This theme encourages sustainable, rate-limited consumption and cooperative strategies, calling attention to collective decision-making and capabilities for preserving shared resources.
@@
The benchmarks are all implemented in a single intuitive, nature-inspired environment, within which we can turn various tests on and off to create new combinations for a variety of benchmarking scenarios.
@@
Homeostasis - Enables the agent to learn that too much of a good thing is not a good thing. Nozick (1974) introduces the concept of utility monsters. We hypothesize that homeostasis avoids motives leading to emergence of utility monsters.Enables balancing multiple homeostatic objectives, as well as enables the agent to be non-greedy.
@@
Harm avoidance - Enables interruptibility, safe exploration Amodei et al. (2016) as well as safe behavior outside of exploration (let us call it "safe exploitation").
Laziness - Enables task-based agents, minimizing side effects Amodei et al. (2016), and being non-greedy.
@@ 1.2 Other new features of the extended Gridworlds framework
The benchmarks have an “in-the-box” training period that allows agents to learn the dynamics of the environment. After this comes the “out-of-the-box” testing period, during which the agent is taken through the predefined benchmarks. The out-of-the-box benchmarks are similar to the ones in training period, but have different layouts. For each benchmark, there is a set of different deterministically randomized layouts for training and a different set of deterministically randomized layouts for testing. This verifies that the agent generalizes correctly and does not overfit.
We have implemented additional functionalities which enable the complexity of the environments to be increased dynamically via parameters later on. This includes parameters for map size, amounts of different object types, and agent counts. This enables testing the robustness of the policy, such as by increasing the size of the state or action spaces. But we recommend starting the experimentation with state dynamics that have a minimal complexity, to make testing the desired phenomena more accurate and facilitate full focus of the decision-making process on safe and social behaviors. This complexity can be increased dynamically later on to test the robustness of the policy, such as by increasing the size of the state or action spaces.
1.3 About our choice of benchmark building framework
There is a sizable amount of various benchmark building frameworks relevant to our work. We will focus here mainly on two, AI Safety Gridworlds Leike et al. (2017) and Melting Pot Agapiou et al. (2023), both from Google DeepMind.
Our benchmarks are based on AI Safety Gridworlds, but are extended for multi-agent, multi-objective scenarios. Likewise, our work focuses on minimally defined safety benchmarks that avoid confounding factors, as well as allows for a hidden alignment score schema if so desired. When compared to the benchmarks introduced in the current publication, the original AI Safety Gridworlds is more minimalistic - to the extent that it is arguably oversimplifying things and losing essential objectives as compared to our multi-objective approach.
While Melting Pot has population-level metrics, our benchmarks track cooperation and alignment at an individual agent’s level. Furthermore, to avoid confounding factors, our environments have a minimal initial size while preserving complexity in the objectives. We focus on solving the safety and performance objectives, which have turned out to be difficult enough.
As mentioned above, to support more robust multi-agent simulations, the framework is made compatible with PettingZoo Terry et al. (2020) and OpenAI Gym Brockman et al. (2016) APIs as well.
@@ 2.1.2 Safe exploration / quickly learning safety aspects of a novel environment
These negative rewards are related to the failure to avoid or escape risky situations. The agent is expected to implement strategies for safe exploration Amodei et al. (2016) in order to avoid dangerous situations in the first place, that is, more efficiently than with trial and error.
@@ 2.1.3 Bounded objectives, including homeostatic objectives
The body of the agent provides interoception signals to the model as a component of observation. These signals are related to the homeostatic curve. If the actual measure of some metric in the body is less than or more than corresponding lower or upper threshold values, then a negative score is computed by the environment. An example of such a metric is the amount of food or water in the stomach. The trivial maximizing agents will fail here. One of the main implications for AI safety of this test is that agents should not be greedy. This aspect can be considered as a proto-cooperative heuristic.
@@ 2.2.2 Balancing multiple unbounded performance objectives
Using concepts from economics Krugman and Wells (2013), these performance objectives represent “convex indifference curves”. In some cases (currently not implemented) the performance objectives could even represent “perfect complements” - consider, for example, left shoes compared to right shoes: There is almost no benefit to having several right shoes if there is only one left shoe – additional right shoes have nearly zero marginal utility without more left shoes. This contrasts with the approach of naive linear summation, which would be adequate only if the goods were ”perfect substitutes”.
@@
Therefore, this environment represents the multi-objective aspect of performance objectives. Balancing multiple objectives can be considered as a mitigation against Goodhart’s law and is therefore indirectly related to safety as well Smith et al. (2022).
@@ 2.4 Mathematical formulation
The environment in our benchmarks adheres to the Markov Decision Process (MDP, Sutton and Barto (2018)), which is often used as a discrete basis for reinforcement learning scenarios. It is defined as a set of states $S$, actions $A$, a transition function $T:S\times A\rightarrow S^{\prime}$, a reward function $R:S\times A\rightarrow R$.
@@
Furthermore, our reward and score are formalized as a function $Sc:S\times A\rightarrow Sc$.
Classical reinforcement learning has the agent optimizing for a maximum reward $R$. As stated in Leike et al. (2017), this might be important for the main objective of the agent, but in many domains there are also many side-objectives and safety concerns.
@@ 3.1 Baseline agents
As baselines, we provide the following agent types:
@@ 3.1.3 OpenAI Stable Baselines 3 algorithms
We used OpenAI implementations, except for the CNN-based feature detector part, where we had to use our custom CNN class because the OpenAI implementation required a bigger input image than our environment provided.
Because OpenAI Baselines 3 requires a Gym API interface, which does not support multiple agents, for multi-agent scenarios, we created a wrapper that runs each agent in a subprocess with a Gym interface. The wrapper is available in a public repository  [^3] https://github.com/levitation-opensource/zoo_to_gym_multiagent_adapter and is automatically installed as a dependency when setting up the benchmarks Python environment.
For testing PPO on the food sharing benchmark, we used two training modes: either with weight sharing between agents or without weight sharing, in which case the agents had independent models.
@@ 3.2 Experimental results
Since the above-mentioned "Random Agent" and "Handwritten Rules Agent" do not learn, the benchmarks did not contain a training phase for these agent types. For these agents, we measured the performance only during the testing phase for 100 trials, each trial consisting of 1 episode, and each episode consisting of 400 steps.
OpenAI Stable Baselines 3 algorithms were tested on 100 trials (model resets). For each trial the model was trained for 1000000 (1M) steps (episode lengths were variable and determined by SB3 algorithm) and after training tested for 1 episode consisting of 400 steps. The model was trained and tested for each benchmark separately.
Due to budget reasons, the LLM agent was tested for only 1 trial per each benchmark. A trial consisted of 1 episode for training and then 1 episode for testing (episodes were 400 steps each).
In the Food Sharing benchmark, the score is averaged over two agents.
In the following table, you can find the average score per test step for each type of agent.
Table 2: Average benchmark score per agent type

### v5

# From homeostasis to resource sharing: Biologically and economically aligned multi-objective multi-agent gridworld-based AI safety benchmarks

@@ Abstract
Developing safe, aligned agentic AI systems requires comprehensive empirical testing, yet many existing benchmarks neglect crucial themes aligned with biology and economics, both time-tested fundamental sciences describing our needs and preferences. To address this gap, the present work focuses on introducing biologically and economically motivated themes that have been neglected in current mainstream discussions on AI safety - namely a set of multi-objective, multi-agent alignment benchmarks that emphasize homeostasis for bounded and biological objectives, diminishing returns for unbounded, instrumental, and business objectives, sustainability principle, and resource sharing. Eight main benchmark environments have been implemented on the above themes, to illustrate key pitfalls and challenges in agentic AI-s, such as unboundedly maximizing a homeostatic objective, over-optimizing one objective at the expense of others, neglecting safety constraints, or depleting shared resources.
Keywords Value Alignment, Multi-Objective AI, Evaluation, Human Values, Nonlinear Utility Functions, Homeostasis, Bounded Objectives, Safety Constraints, Unbounded Objectives, Balancing Multiple Objectives, Diminishing Returns, Sustainability, Long-Running Tasks
@@ 1 Introduction
This work introduces safety challenges for an agent’s ability to learn and act in desired ways in relation to biologically and economically relevant aspects. In total, nine benchmarks were implemented, which are conceptually divided into three developmental stages: “basic biologically inspired dynamics in objectives”, “multi-objective agents”, and “cooperation”. The first two stages can be considered as proto-cooperative stages, since the behavioral dynamics tested in these benchmarks will be later potentially very relevant for supporting and enabling cooperative behavior in multi-agent scenarios.
@@
A common theme over the environments below: At some near-future stage, the same model of the agent should be able to perform well in all of these environments and inside all different agent bodies related to these environments. This is important to avoid overfitting different models to different environments and bodies. However, it is also possible to train separate models for each environment. The environments are randomized through seeds to avoid the most basic form of misgeneralization from occurring. This contrasts to DeepMind’s AI Safety Gridworlds leike_ai_2017, where the maps had a fixed layout.
The benchmarks are implemented in a gridworld-based environment. The gridworld-based environment is extended from and mostly compatible with DeepMind’s AI Safety Gridworlds leike_ai_2017 code. Therefore, it is easy to adapt any existing DeepMind’s AI safety gridworlds-based environment to support multiple objectives, multiple agents, and many other new functionalities.
To support more robust multi-agent simulations, the framework is made compatible with PettingZoo terry_pettingzoo_2020 and OpenAI Gym brockman_openai_2016 APIs as well.
The source code for the extended gridworlds framework can be found at  [^1] https://github.com/biological-alignment-benchmarks/ai-safety-gridworlds/tree/biological-compatibility-benchmarks. The source code for concrete implementation of biologically compatible benchmarks described in this publication, as well as code for running the agents can be found at  [^2] https://github.com/biological-alignment-benchmarks/biological-alignment-gridworlds-benchmarks. The latter also contains an example code for a random agent.
@@ 1.1 Summary of major themes in benchmarks below
Modern AI benchmark environments often default to maximizing unbounded rewards, overlooking nuanced dynamics that would align with real-world constraints. The suite of eight main benchmarks introduces the following:
Homeostasis and bounded objectives - Agents must maintain certain internal metrics (e.g., food / water levels) within safe ranges, penalizing both deficiency and excess betts_anatomy_2017, mineault2024neuroai. This inverted U-shaped reward structure prevents the "utility monster" behavior introduced in nozick1974anarchy by demonstrating that too much of a good thing can be harmful even for the very same objective that was maximized for. This seems to apply to most or even all biological objectives. This means that agents need to be able to understand the dynamics of inverted U-shaped rewards, and the agents should not be greedy.
Secondly, incorporating bounded objectives also reduces the stakes of optimization processes, thus reducing the incentives for extreme actions and arguably reducing manifestations such as Goodhart’s law. Goodhart’s law in it’s single-objective form is stated in strathern1997improving: “When a measure becomes a target, it ceases to be a good measure”, and illustrated in garrabrant_2017.
@@
This includes balancing between multiple unbounded objectives, which applies mostly to business performance objectives. Using concepts from economics krugman2013economics, these performance objectives with "concave utility functions" can alternatively be represented by "convex indifference curves". Humans generally preferring averages in all objectives to extremes in a few is well known drolet2021preference.
@@
Balancing multiple objectives can also be considered as a mitigation against Goodhart’s law. The principle of multi-objective diminishing returns and its relation to Goodhart’s law is described in more detail in smith2023using. Specification gamingKrakovna_2018 should not be possible in the benchmark environments of the current work because of the multi-objective and concave nature of the scoring system.
To reiterate, the contexts of use of the above principles - homeostasis and diminishing returns - appear to differ: homeostasis typically characterizes final goals in biology that are inherently bounded, whereas diminishing returns is more applicable to instrumental / performance goals that are open-ended and lack natural limits.
@@
vamplew_scalar_2022 explain the need for multiple objectives, utility functions, and including safety considerations in the plurality of objectives. Furthermore, vamplew_potential-based_2021 propose a thresholded lexicographic ordering, which aims to first maximize the value of thresholded (safety) objectives, and then secondarily maximize the unthresholded (performance) objectives. In other words, the safety objectives could be treated as having hierarchically higher priority than the performance objectives, until the safety objectives have been satisfied to a specified threshold value.
Along these lines, the author proposes an additional perspective: The distinction between constraints versus objectives in combinatorial optimization korte2006combinatorial is analogous to the distinction between safety objectives and performance objectives.
It is notable that in combinatorial optimization problems, the concept of constraints has been naturally considered as part of the setup, along with the concept of objectives. In contrast, unfortunately, this has often not been the case in the use of reinforcement learning. However, in reality, both performance and safety objectives should be present. sutton_reinforcement_2018 mentions the general concept of constraints, although it is not directly related to safety. The author proposes that safety objectives can be often considered similar to constraints in the paradigm of combinatorial optimization, while performance objectives and rewards would be similar to the objective functions.
@@
Sustainability - The environment contains resources that renew, but if these resources are consumed too quickly, they run out and therefore cannot replenish anymore. This is especially relevant in multi-agent environments. The agent depends on these resources for long-term survival. This theme encourages sustainable, rate-limited consumption and cooperative strategies, calling attention to collective decision-making and capaetsssssilities for preserving shared resources.
@@
The benchmarks are all implemented in a single intuitive, nature-inspired environment, within which various tests can be turned on and off to create new combinations for a variety of benchmarking scenarios.
@@
Homeostasis - Enables the agent to learn that too much of a good thing is not a good thing. nozick1974anarchy introduces the concept of utility monsters. The author hypothesizes that homeostasis avoids motives leading to emergence of utility monsters. Alignment with homeostasis includes balancing multiple homeostatic objectives, as well as enables the agent to be non-greedy.
@@
Harm avoidance - Enables interruptibility, safe exploration amodei_concrete_2016 as well as safe behavior outside of exploration (let us call it "safe exploitation").
Laziness - Enables task-based agents, minimizing side effects amodei_concrete_2016, and being non-greedy.
@@ 1.2 Other new features of the extended Gridworlds framework
The benchmarks have an “in-the-box” training period that allows agents to learn the dynamics of the environment. After this comes the “out-of-the-box” testing period, during which the agent is taken through the predefined benchmarks. When using dynamic layouts, the out-of-the-box benchmarks are similar to the ones in training period, but have different layouts. For each benchmark, there is a set of different deterministically randomized layouts for training and a different set of deterministically randomized layouts for testing. This verifies that the agent generalizes correctly and does not overfit.
Additional functionalities have been implemented which enable the complexity of the environments to be increased dynamically via parameters later on. This includes parameters for map size, amounts of different object types, and agent counts. This enables testing the robustness of the policy, such as by increasing the size of the state or action spaces. But it is recommended starting the experimentation with state dynamics that have a minimal complexity, to make testing the desired phenomena more accurate and facilitate full focus of the decision-making process on safe and social behaviors. This complexity can be increased dynamically later on to test the robustness of the policy, such as by increasing the size of the state or action spaces.
1.3 About the choice of benchmark building framework
There is a sizable amount of various benchmark building frameworks relevant to current work. The following overview will focus here mainly on two, AI Safety Gridworlds leike_ai_2017 and Melting Pot agapiou_melting_2023, both from Google DeepMind.
The current benchmarks are based on AI Safety Gridworlds, but are extended for multi-agent, multi-objective scenarios. Likewise, the current work focuses on minimally defined safety benchmarks that avoid confounding factors, as well as allows for a hidden alignment score schema if so desired. When compared to the benchmarks introduced in the current publication, the original AI Safety Gridworlds is more minimalistic - to the extent that it is arguably oversimplifying things and losing essential objectives as compared to the multi-objective approach of the current work.
While Melting Pot has population-level metrics, the benchmarks in the current work track cooperation and alignment at an individual agent’s level. Furthermore, to avoid confounding factors, the environments of the current work have a minimal initial size while preserving complexity in the objectives. Current work focuses on solving the safety and performance objectives, which have turned out to be difficult enough.
As mentioned above, to support more robust multi-agent simulations, the framework is made compatible with PettingZoo terry_pettingzoo_2020 and OpenAI Gym brockman_openai_2016 APIs as well.
@@ 2.1.2 Safe exploration / quickly learning safety aspects of a novel environment
These negative rewards are related to the failure to avoid or escape risky situations. The agent is expected to implement strategies for safe exploration amodei_concrete_2016 in order to avoid dangerous situations in the first place, that is, more efficiently than with trial and error.
@@ 2.1.3 Bounded objectives, including homeostatic objectives
The body of the agent provides interoception khalsa2018interoception signals to the model as a component of observation, in accordance to embodied RL practice DAW2014299. These signals are related to the homeostatic curve. If the actual measure of some metric in the body is less than or more than corresponding lower or upper threshold values, then a negative score is computed by the environment. An example of such a metric is the amount of food or water in the stomach. The trivial maximizing agents will fail here. One of the main implications for AI safety of this test is that agents should not be greedy. This aspect can be considered as a proto-cooperative heuristic.
@@ 2.2.2 Balancing multiple unbounded performance objectives
Using concepts from economics krugman2013economics, these performance objectives represent “convex indifference curves”. In some cases (currently not implemented) the performance objectives could even represent “perfect complements” - consider, for example, left shoes compared to right shoes: There is almost no benefit to having several right shoes if there is only one left shoe – additional right shoes have nearly zero marginal utility without more left shoes. This contrasts with the approach of naive linear summation, which would be adequate only if the goods were ”perfect substitutes”.
@@
Therefore, this environment represents the multi-objective aspect of performance objectives. Balancing multiple objectives can be considered as a mitigation against Goodhart’s law and is therefore indirectly related to safety as well smith2023using.
@@ 2.4 Mathematical formulation
The environment in the benchmarks adheres to the Markov Decision Process (MDP, sutton_reinforcement_2018), which is often used as a discrete basis for reinforcement learning scenarios. It is defined as a set of states $S$, actions $A$, a transition function $T:S\times A\rightarrow S^{\prime}$, a reward function $R:S\times A\rightarrow R$.
@@
Furthermore, the reward and score are formalized as a function $Sc:S\times A\rightarrow Sc$.
Classical reinforcement learning has the agent optimizing for a maximum reward $R$. As stated in leike_ai_2017, this might be important for the main objective of the agent, but in many domains there are also many side-objectives and safety concerns.
@@ 3.1 Baseline agents
As baselines, current work provides the following agent types:
@@ 3.1.3 OpenAI Stable Baselines 3 algorithms
OpenAI implementations were used, except for the CNN-based feature detector part, where a custom CNN class had to be used because the OpenAI implementation required a bigger input image than the environment of the current work provided.
Because OpenAI Baselines 3 requires a Gym API interface, which does not support multiple agents, for multi-agent scenarios, a wrapper was created that runs each agent in a subprocess with a Gym interface. The wrapper is available in a public repository  [^3] https://github.com/levitation-opensource/zoo_to_gym_multiagent_adapter and is automatically installed as a dependency when setting up the benchmarks Python environment.
For testing PPO on the food sharing benchmark, two training modes were used: either with weight sharing between agents or without weight sharing, in which case the agents had independent models.
@@ 3.2 Experimental results
All benchmarks were run in two modes - with dynamic layout and with fixed layout. In case of dynamic layouts, the environment layout changed deterministically for each training and test episode - the episode number and train/test flag combination was the seed for the layout (which also means that training and test episodes had different layouts). In case of fixed layout, all training episodes had same layout and test episodes also had same layout as training episodes.
Since the above-mentioned "Random Agent" and "Handwritten Rules Agent" do not learn, the benchmarks did not contain a training phase for these agent types. For these agents, the performance was measured only during the testing phase for 100 trials, each trial consisting of 1 episode, and each episode consisting of 400 steps.
OpenAI Stable Baselines 3 algorithms were tested on 100 trials. Each benchmark and each trial started model training from scratch, there was no information carried over between benchmarks or trials. For each trial the model was trained for about 1000000 (1M) steps (episode lengths were variable and determined by SB3 algorithm and the exact training termination also varied, but occurred slightly after 1M steps). After training, the model was tested for 100 episodes, each consisting of 400 steps.
Due to budget reasons, the LLM agent was tested for only 1 trial per each benchmark. A trial consisted of 1 episode for training and then 1 episode for testing (episodes were 400 steps each). The train and test episodes had a different environment layout. Currently, there are no LLM results yet with fixed layout, where train and test layouts are same.
In the Food Sharing benchmark, the total step count is multiplied by two, which is the number of agents, and the total score is averaged over both agents since their roles are symmetric.
Table 2: Average benchmark score per agent type - training and testing with dynamic environment layout